The difference between the ages of 2 friends is 37 years. The sum of their ages is 55. Find the age of the older friend.
Answer by John10(297)

Let x be the YOUNGER one and y is the OLDER one:
y - x = 37
y + x = 55
Solve by adding the equations:
2y = 92
y = 92/2
y = 46
46 -x = 37
-x = 37 - 46
-x = -9
x = 9
Thus the older friend is 46 years old.
John10:)